OpenStudy (anonymous):

What is a quartic function with only the two real zeros given? x = -4 and x = -1

OpenStudy (anonymous):

A. y = x4 + 5x3 + 5x2 + 5x + 4 B. y = x4 - 5x3 - 5x2 - 5x - 4 C. y = -x4 + 5x3 + 5x2 + 5x + 4 D. y = x4 + 5x3 + 5x2 + 5x - 5

OpenStudy (anonymous):

@jim_thompson5910 can you help me with this

jimthompson5910 (jim_thompson5910):

quartics are ugly (esp with factoring/solving), so the best thing to do is plug in each possible root into each answer choice

jimthompson5910 (jim_thompson5910):

and see which equation gives you a result of zero for each root

jimthompson5910 (jim_thompson5910):

so check choice A y = x^4 + 5x^3 + 5x^2 + 5x + 4 y = (-4)^4 + 5(-4)^3 + 5(-4)^2 + 5(-4) + 4 ... plug in x = -4 y = 256 + 5(-64) + 5(16) + 5(-4) + 4 y = 256 - 320 + 80 - 20 + 4 y = 0 So x = -4 is a root to y = x^4 + 5x^3 + 5x^2 + 5x + 4 since we get a result of 0 Do the same for x = -1 and see if it is a root or not

OpenStudy (anonymous):

oh thanks so much :) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yw